OpenStudy (anonymous):

true or false, you can use the law of sines if you are given a=31, b=52 and angle B=56

sam (.sam.):

True \[\frac{\sin(56)}{52}=\frac{\sin(a)}{31}\]

OpenStudy (anonymous):

thank you